# Training-free Zero-shot Composed Image Retrieval via Weighted Modality Fusion and Similarity

### v3

@@ Abstract
Composed image retrieval (CIR), which formulates the query as a combination of a reference image and modified text, has emerged as a new form of image search due to its enhanced ability to capture users’ intentions. However, training a CIR model in a supervised manner typically requires labor-intensive collection of (reference image, text modifier, target image) triplets. While existing zero-shot CIR (ZS-CIR) methods eliminate the need for training on specific downstream datasets, they still require additional pretraining on large-scale image datasets. In this paper, we introduce a training-free approach for ZS-CIR. Our approach, Weighted Modality fusion and similarity for CIR (WeiMoCIR), operates under the assumption that image and text modalities can be effectively combined using a simple weighted average. This allows the query representation to be constructed directly from the reference image and text modifier. To further enhance retrieval performance, we employ multimodal large language models (MLLMs) to generate image captions for the database images and incorporate these textual captions into the similarity computation by combining them with image information using a weighted average. Our approach is simple, easy to implement, and its effectiveness is validated through experiments on the FashionIQ and CIRR datasets.
@@ 1 Introduction
Given a reference image and a textual description, composed image retrieval (CIR) is to find a target image, that visually resembles the reference image and also incorporates modifications specified by the text, from an image database. Compared to using a single modality, either images or texts, which could have limited expressiveness in describing users’ preferences, CIR allows users to better express their preferences with multimodal queries. Consequently, recent years have witnessed significant research efforts in CIR [2, 3, 14, 17].
Multi-modality queries pose a unique challenge to image retrieval because the retrieval algorithms need to understand what to retain in the visual content and what to modify based on finer details expressed in the natural language. Previous studies [3, 13] have tackled this by training CIR models on large-scale (reference image, text modifier, target image) triplets to learn how to fuse the reference image and the text modifier into a query representation for retrieval. However, collecting such triplets requires significant human effort, and these supervised approaches usually exhibit limited performance on unseen data due to being tuned on specific datasets. As a result, zero-shot CIR (ZS-CIR), which is able to perform CIR without the need for supervised learning on downstream data, has gained attention. However, existing ZS-CIR approaches, such as Pic2Word [17] and SEARLE [2], still require training on large image datasets to learn mappings for converting images into pseudo-word tokens, which are then combined with textual descriptions for text-to-image retrieval. This additional training can be resource-intensive.
In this paper, we introduce a training-free approach for ZS-CIR. Our approach leverages existing vision-language models (VLMs) and multimodal large language models (MLLMs). Specifically, we use a VLM, such as CLIP [15], to obtain visual features from the reference image and textual features from the text modifier. Given that images and texts are aligned within a shared space in the VLMs, we assume that the fused query representation can be obtained through a weighted combination of the visual and textual representations. During retrieval, unlike previous methods that rely solely on comparing the query feature and database image features for retrieval, our method also incorporates the textual information of the target images for similarity computation. This is achieved by using an MLLM, such as Gemini [7], to generate captions for each database image. To capture different aspects of the image content, multiple descriptions are generated for each database image. By incorporating these generated descriptions, our approach considers both the similarities between the query and visual features of the database images and the similarities between the query and textual features of database images through a weighted average. We term our approach WeiMoCIR, which stands for Weighted Modality fusion and similarity for CIR.
@@ 2 Related Work
In CIR, the query consists of two modalities: a reference image and textual modifications, with goal of retrieving images that meet the conditions specified in the query. Earlier CIR methods [3, 6, 21] have focused on developing multimodal fusion techniques to combine the visual features and textual features to obtain the query representation. For instance, CLIP4CIR [3] trains a combiner to fuse the information from two modalities. However, these methods require triplets for specific downstream applications for model training, and collecting such triplets is labor-intensive. As a result, recent research has explored ZS-CIR. Representative methods, like Pic2Word [17] and SEARLE [2], learn to project images into the textual space using large-scale image datasets. With the learned mappings from images to texts, the projected image is then combined with the text modifier for text-to-image retrieval without needing to train on downstream datasets. CIReVL [10] further eliminates the need for additional training to learn the image-to-text mappings and achieves training-free ZS-CIR. It utilizes a pretrained VLM [11, 15] to generate a textual description for the reference image. This description and the text modifier are fed into an LLM [4, 16] to produce a textual description of the desired image for text-to-image retrieval.
Similar to CIReVL, our approach is also training-free, but differs in two key ways: (1) our method composes the query representation from both visual and text representation while CIReVL translates images to texts; and (2) our method utilizes an MLLM to generate captions for the database images and considers both visual and textual information during retrieval whereas CIReVL considers only the visual features of the database images.
@@ 4.1 Experimental Benchmarks and Protocols
The evaluation metric used is the top K retrieval results (R@K). For FashionIQ, we compute R@K for each of the three categories and their average. Since the test set is no longer available for evaluation, we follow previous studies [2, 8, 9, 10, 17, 23] and report results on the validation set. For CIRR, we evaluate full-set ranking results using R@1, R@5, R@10, and R@50 metrics. For the subset group, we assess performance at R@1, R@2, and R@3. We use the validation set to select hyperparameters $\alpha$ and $\beta$ and report the results on the Test1 set.
@@ 4.2 Implementation details
Table 1: R@K results on Fashion-IQ. The best scores for each backbone are highlighted in bold. The pretrained CLIP is used as the feature extractor.
Shirt Dress Toptee Average Backbone Method R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50 Methods that need an additional pre-traning step ViT L/14 Pic2Word [17] 26.20 43.60 20.00 40.20 27.90 47.40 24.70 43.70 SEARLE-XL [2] 26.89 45.58 20.48 43.13 29.32 49.97 25.56 46.23 SEARLE-XL-OTI [2] 30.37 47.49 21.57 44.47 30.90 51.76 27.61 47.90 CompoDiff [8] 37.69 49.08 32.24 46.27 38.12 50.57 36.02 48.64 LinCIR [9] 29.10 46.81 20.92 42.44 28.81 50.18 26.28 46.49 MagicLens [23] 32.70 53.80 25.50 46.10 34.00 57.70 30.70 52.50 ViT H/14 Pic2Word [9] 36.90 55.99 28.01 51.51 40.18 62.01 35.03 56.50 SEARLE [9] 36.46 55.45 28.46 51.07 38.81 60.89 34.57 55.80 LinCIR [9] 36.90 57.75 29.80 52.11 42.07 62.52 36.26 57.46 ViT G/14 Pic2Word [9] 33.17 50.39 25.43 47.65 35.24 57.62 31.28 51.89 SEARLE [9] 36.46 55.35 28.16 50.32 39.83 61.45 34.81 55.71 CompoDiff [8] 41.31 55.17 37.78 49.10 44.26 56.41 39.02 51.71 LinCIR [9] 46.76 65.11 38.08 60.88 50.48 71.09 45.11 65.69 Training-free methods ViT L/14 CIReVL [10] 29.49 47.40 24.79 44.76 31.36 53.65 28.55 48.57 WeiMoCIR (Ours) 32.78 48.97 25.88 47.29 35.95 56.71 31.54 50.99 ViT H/14 WeiMoCIR (Ours) 36.56 53.58 28.76 48.98 39.72 59.87 35.01 54.14 ViT G/14 CIReVL [10] 33.71 51.42 27.07 49.53 35.80 56.14 32.19 52.36 WeiMoCIR (Ours) 37.73 56.18 30.99 52.45 42.38 63.23 37.03 57.29
@@
During retrieval, as the captions of the database images are incorporated into the similarity computation, their quality could impact performance. Our experiments show that when the generated captions are more closely aligned with the text modifiers, our method produces more favorable results. For FashionIQ, the MLLM is instructed to describe the characteristics of the clothing while explicitly avoiding any mention of the person wearing the clothes. For CIRR, we employ a role-playing instruction to guide the MLLM to provide concise descriptions and exclude any descriptions unread to the images. For both datasets, we provide a defined output format and a set of example descriptions, and specify that the output should consist of $R$ captions, with $R$ set to 3.
@@ 4.3 Main Experimental Results
Table 1 shows the comparison of our approach with other ZS-CIR methods, including those requiring additional pretraining and those that are training-free. It is observed that zero-shot methods with additional pretraining, such as MagicLens and LinCIR, achieve the best performance. This superior performance is primarily because the additional pretraining adapts the model to be more suitable for the CIR task.
It is worth noting that training-free approaches sometimes outperform some methods that involve additional training, demonstrating that favorable results can be achieved by leveraging pretrained models and a well-designed retrieval pipeline. Our method consistently surpasses the other training-free approach, CIReVL, across all metrics. Furthermore, our results improve with the use of larger CLIP models on this dataset. These results validate that leveraging pretrained VLMs and MLLMs, employing a simple weighted sum fusion of the reference image and text features, and incorporating textual information into similarity computation are effective strategies for CIR.
@@
Table 2 shows the results. Similar to the findings on FashionIQ, methods requring additional pretraining, particularly those utilizing larger CLIP models, generally have an advantage over training-free methods. Our approach yields results comparable to CIReVL when using CLIP ViT L/14 but underperforms with CLIP ViT G/14. However, unlike the results on FashionIQ, our method does not consistently yield better results with larger models. This could be due to two reasons: (1) Although larger models generally have stronger generalization capabilities, their performance on specific datasets may not always exceed that of smaller models. (2) Our approach relies heavily on pretrained VLMs for feature extraction, and the choice of VLM can significantly influence retrieval performance, as confirmed by the ablation study presented next.
@@
Fine-tuned on Flickr30k   Fine-tuned on COCO
Figure 2: Effects of $\alpha$ and $\beta$ on retrieval performance on FashionIQ. Left: average R@50 using CLIP with ViT G/14. Right: average R@50 using BLIP with ViT-B.
@@ 4.4 Ablation Study
In Table 3, we study the impact of pretrained VLMs on retrieval performance. The effectiveness of CLIP versus BLIP varies by dataset: CLIP outperforms BLIP on FashionIQ, while both models exhibit comparable performance on CIRR. When using the BLIP models that have been further fine-tuned on COCO or Flickr30k, our method achieves significantly better results on both datasets compared to using the BLIP model without additional fine-tuning. This improvement is likely due to the high relevance of COCO and Flickr30k images to FashionIQ and CIRR images. These findings suggest a strong correlation between the VLMs’ pretraining strategies and datasets and their performance on specific downstream datasets. As our method is training-free, it offers the flexibility to easily employ different VLMs depending on the datasets and applications.
(a) Retrieval results of example queries from Fashion-IQ.
(b) Retrieval results of example queries from CIRR.
Figure 3: Retrieval results of example queries from FashionIQ and CIRR. Our method successfully retrieves the desired images, highlighted with a green box, demonstrating its ability to perform CIR for a wide variety of text modifiers.
@@
As observed, when only a single image caption is included in the similarity computation, the first caption generated by the MLLM yields better results than the other two, suggesting it better captures the image content. Incorporating multiple captions further enhances retrieval performance, mainly because combining more captions provides a more complete description of the image. We emphasize that the performance gains from including multiple captions is achieved by guiding MLLMs with well-designed prompts, without any additional human annotations.
@@ 4.5 Qualitative Results
We show the CIR results produced our method in Figure 3. These examples clearly show that our training-free ZS-CIR method can effectively retrieve desired images for a wide range of text modifiers.
@@ 5 Conclusion
We have presented a training-free approach for ZS-CIR. By leveraging pretrained VLMs and MLLMs, our approach integrates information from different modalities via a weighted average. This results in an easy approach to construct the query representation and allows for the incorporation of textual information into the similarity computation. Our approach performs CIR without any training and hence has the flexibility to use different VLMs and MLLMs depending on the datasets and applications. Experiments and analyses on two datasets demonstrate the effectiveness of our method.

### v4

@@ Abstract
Composed image retrieval (CIR), which formulates the query as a combination of a reference image and modified text, has emerged as a new form of image search due to its enhanced ability to capture user intent. However, training a CIR model in a supervised manner typically requires labor-intensive collection of (reference image, text modifier, target image) triplets. While existing zero-shot CIR (ZS-CIR) methods eliminate the need for training on specific downstream datasets, they still require additional pretraining on large-scale image datasets. In this paper, we introduce a training-free approach for ZS-CIR. Our approach, Weighted Modality fusion and similarity for CIR (WeiMoCIR), operates under the assumption that image and text modalities can be effectively combined using a simple weighted average. This allows the query representation to be constructed directly from the reference image and text modifier. To further enhance retrieval performance, we employ multimodal large language models (MLLMs) to generate image captions for the database images and incorporate these textual captions into the similarity computation by combining them with image information using a weighted average. Our approach is simple, easy to implement, and its effectiveness is validated through experiments on the FashionIQ and CIRR datasets. Code is available at https://github.com/whats2000/WeiMoCIR.
@@ 1 Introduction
Given a reference image and a textual description, composed image retrieval (CIR) is to find a target image, that visually resembles the reference image and also incorporates modifications specified by the text, from an image database. Compared to using a single modality (e.g., images or texts alone), which could have limited expressiveness in capturing users’ preferences, CIR allows users to better express their preferences with multimodal queries. Consequently, recent years have witnessed significant research efforts in CIR [2, 3, 14, 17].
Multimodal queries pose a unique challenge to image retrieval because the retrieval algorithms need to understand what to retain in the visual content and what to modify based on finer details expressed in the natural language. Previous studies [3, 13] have tackled this by training CIR models on large-scale (reference image, text modifier, target image) triplets to learn how to fuse the reference image and the text modifier into a query representation for retrieval. However, collecting such triplets requires substantial human effort, and the supervised approaches usually exhibit limited performance on unseen data due to being tuned on specific datasets. Consequently, zero-shot CIR (ZS-CIR), capable of performing CIR without supervised learning on downstream data, has gained attention. However, existing ZS-CIR methods, such as Pic2Word [17] and SEARLE [2], still require training on large image datasets to learn mappings for converting images into pseudo-word tokens, which are then combined with textual descriptions for text-to-image retrieval. This additional training can be resource-intensive.
In this paper, we introduce a training-free approach for ZS-CIR. Our approach leverages existing vision-language models (VLMs) and multimodal large language models (MLLMs). Specifically, we use a VLM, such as CLIP [15], to obtain visual features from the reference image and textual features from the text modifier. Given that images and texts are aligned within a shared space in the VLMs, we assume that the fused query representation can be obtained through a weighted combination of the visual and textual representations. During retrieval, unlike previous methods that rely solely on comparing the query feature and database image features for retrieval, our method also incorporates the textual information of the database images for similarity computation. This is achieved by using an MLLM, such as Gemini [7], to generate captions for each database image. To capture different aspects of the image content, multiple descriptions are generated for each database image. By incorporating these generated descriptions, our approach considers both the similarities between the query and visual features of the database images and the similarities between the query and textual features of database images through a weighted average. We term our approach WeiMoCIR, which stands for Weighted Modality fusion and similarity for CIR.
@@ 2 Related Work
In CIR, the query comprises two modalities—a reference image and a text modifier—with the goal of retrieving images that satisfy the conditions specified in the query. Earlier CIR methods [3, 6, 21] have focused on developing multimodal fusion techniques to combine the visual features and textual features to obtain the query representation. For instance, CLIP4CIR [3] trains a combiner to fuse the information from two modalities. However, these methods require triplets for specific downstream applications for model training, and collecting such triplets is labor-intensive. As a result, recent research has explored ZS-CIR. Representative methods, like Pic2Word [17] and SEARLE [2], learn to project images into the textual space using large-scale image datasets. With the learned mappings from images to texts, the projected pseudo-word tokens are then combined with the text modifier for text-to-image retrieval without needing to train on downstream datasets. CIReVL [10] further eliminates the need for additional training to learn the image-to-text mappings and achieves training-free ZS-CIR. It utilizes a pretrained VLM [11, 15] to generate a textual description for the reference image. This description and the text modifier are fed into an LLM [4, 16] to produce a textual description of the desired image for text-to-image retrieval. LDRE [23] leverages a pretrained image captioner to generate multiple descriptions for the reference images and employs an LLM to reason based on these descriptions and the modified text. This approach demonstrates that multiple image captions can capture the diverse semantics of the composed result, thereby enhancing performance.
Similar to CIReVL and LDRE, our approach is training-free but differs in two key aspects: (1) our method composes the query representation from both visual and text representation while CIReVL and LDRE translate images to texts; and (2) our method utilizes an MLLM to generate captions for the database images and considers both visual and textual information during retrieval. In contrast, CIReVL and LDRE only consider the visual features of database images.
@@ 4.1 Experimental Benchmarks and Protocols
The evaluation metric is the top K retrieval results (R@K). For FashionIQ, we compute R@K for each of the three categories and their average. Since the test set is no longer available for evaluation, we follow previous studies [2, 8, 9, 10, 17, 24] and report results on the validation set. For CIRR, we evaluate full-set ranking results using R@1, R@5, R@10, and R@50 metrics. For the subset group, we assess performance using R@1, R@2, and R@3. We use the validation set to select hyperparameters $\alpha$ and $\beta$ and report the results on the Test1 set.
@@ 4.2 Implementation details
Table 1: R@K results on FashionIQ. The best scores for each backbone are highlighted in bold. The pretrained CLIP is used as the feature extractor.
Shirt Dress Toptee Average Backbone Method R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50 Methods that need an additional pre-traning step ViT L/14 Pic2Word [17] 26.20 43.60 20.00 40.20 27.90 47.40 24.70 43.70 SEARLE-XL-OTI [2] 30.37 47.49 21.57 44.47 30.90 51.76 27.61 47.90 CompoDiff [8] 37.69 49.08 32.24 46.27 38.12 50.57 36.02 48.64 LinCIR [9] 29.10 46.81 20.92 42.44 28.81 50.18 26.28 46.49 MagicLens [24] 32.70 53.80 25.50 46.10 34.00 57.70 30.70 52.50 ViT H/14 Pic2Word [17] 36.90 55.99 28.01 51.51 40.18 62.01 35.03 56.50 SEARLE [2] 36.46 55.45 28.46 51.07 38.81 60.89 34.57 55.80 LinCIR [9] 36.90 57.75 29.80 52.11 42.07 62.52 36.26 57.46 ViT G/14 Pic2Word [17] 33.17 50.39 25.43 47.65 35.24 57.62 31.28 51.89 SEARLE [2] 36.46 55.35 28.16 50.32 39.83 61.45 34.81 55.71 CompoDiff [8] 41.31 55.17 37.78 49.10 44.26 56.41 39.02 51.71 LinCIR [9] 46.76 65.11 38.08 60.88 50.48 71.09 45.11 65.69 Training-free methods ViT B/32 CIReVL [10] 28.36 47.84 25.29 46.36 31.21 53.85 28.29 49.35 LDRE [23] 27.38 46.27 19.97 41.84 27.07 48.78 24.81 45.63 WeiMoCIR (Ours) 29.20 47.15 26.23 46.31 34.17 55.99 29.86 49.82 ViT L/14 CIReVL [10] 29.49 47.40 24.79 44.76 31.36 53.65 28.55 48.57 LDRE [23] 31.04 51.22 22.93 46.76 31.57 53.64 28.51 50.54 WeiMoCIR (Ours) 32.78 48.97 25.88 47.30 35.95 56.71 31.54 50.99 ViT H/14 WeiMoCIR (Ours) 36.56 53.58 28.76 48.98 39.72 59.87 35.01 54.14 ViT G/14 CIReVL [10] 33.71 51.42 27.07 49.53 35.80 56.14 32.19 52.36 LDRE [23] 35.94 58.58 26.11 51.12 35.42 56.67 32.49 55.46 WeiMoCIR (Ours) 37.73 56.18 30.99 52.45 42.38 63.23 37.03 57.29
@@
During retrieval, as the captions of the database images are incorporated into the similarity computation, their quality could impact performance. Our experiments show that when the generated captions are more closely aligned with the text modifiers, our method produces more favorable results. For FashionIQ, the MLLM is instructed to describe the characteristics of the clothes while explicitly avoiding any mention of the person wearing the clothes. For CIRR, we employ a role-playing instruction to guide the MLLM to provide concise descriptions and exclude any descriptions unrelated to the images. For both datasets, we provide a defined output format and a set of example descriptions, and specify that the output should consist of $R$ captions, with $R$ set to 3.
@@ 4.3 Main Experimental Results
Table 1 shows the comparison of our approach with other ZS-CIR methods, including those requiring additional pretraining and those that are training-free. It is observed that the zero-shot methods with additional pretraining, such as MagicLens [24] and LinCIR [9], achieve the best performance. This superior performance is primarily due to the additional pretraining that adapts the model to be more suitable for the CIR task.
It is worth noting that the training-free approaches sometimes outperform some methods that involve additional training, demonstrating that favorable results can be achieved by leveraging pretrained models and a well-designed retrieval pipeline. Our method consistently surpasses the other training-free approaches across most of the metrics. Furthermore, our results improve with the use of larger CLIP models on this dataset. These results validate that leveraging pretrained VLMs and MLLMs, employing a simple weighted fusion of the reference image and text features, and incorporating the textual information into similarity computation are effective strategies for CIR.
@@
Table 2 shows the results. Similar to the findings on FashionIQ, the methods requring additional pretraining, particularly those utilizing larger CLIP models, generally have an advantage over the training-free methods. However, unlike the results on FashionIQ, our method does not consistently yield better results with larger models. Hence, as can be seen, our approach yields results comparable to the other training-free methods with smaller CLIP models but underperforms with the larger CLIP ViT-G/14. This could be due to two reasons: (1) although larger models generally have stronger generalization capabilities, their performance on specific datasets may not always exceed that of smaller models, and (2) our approach relies heavily on pretrained VLMs for feature extraction, and the choice of VLM can significantly impact retrieval performance, as confirmed by the ablation study presented next.
@@
Finetuned on Flickr30k   Finetuned on COCO
Figure 2: Effects of $\alpha$ and $\beta$ on retrieval performance on FashionIQ. Left: average R@50 using CLIP ViT G/14. Right: average R@50 using BLIP ViT-B.
@@ 4.4 Ablation Study
In Table 3, we study the impact of pretrained VLMs on retrieval performance. The effectiveness of CLIP versus BLIP varies by dataset: CLIP outperforms BLIP on FashionIQ, while both models exhibit comparable performance on CIRR. When using the BLIP models that have been further finetuned on COCO or Flickr30k, our method achieves significantly better results on both datasets compared to using the BLIP model without additional finetuning. This improvement is likely due to the high relevance of COCO and Flickr30k images to FashionIQ and CIRR images. These findings suggest a strong correlation between the VLMs’ pretraining strategies and datasets and their performance on specific downstream datasets. As our method is training-free, it offers the flexibility to easily employ different VLMs depending on the datasets and applications.
@@
(a) Retrieval results of example queries from FashionIQ.
(b) Retrieval results of example queries from CIRR.
Figure 3: Retrieval results of example queries from FashionIQ and CIRR. Our method with CLIP ViT L/14 successfully retrieves the desired images, highlighted with green boxes, demonstrating its ability to perform CIR for a wide variety of text modifiers.
@@
As observed, when only a single image caption is included in the similarity computation, the first caption generated by the MLLM yields better results than the other two, suggesting it captures the image content more effectively. Incorporating multiple captions further enhances retrieval performance, mainly because combining more captions provides a more complete description of the image. We emphasize that the performance gains from including multiple captions is achieved by guiding MLLMs with well-designed prompts, without any additional human annotations.
(a) Failure cases of example queries from FashionIQ.
(b) Failure cases of example queries from CIRR.
Figure 4: Failure cases of our method with CLIP ViT L/14 on example queries from FashionIQ and CIRR. Target images are highlighted with green boxes in the last column. As observed, the failures stem from difficulties in capturing more abstract concepts like ‘simple print’ and recognizing specific ‘dog breeds’.
@@ 4.5 Qualitative Results
Figure 3 shows the CIR results from our method on example queries from FashionIQ and CIRR. The text modifiers cover a range of modifications, from color and style adjustments in fashion images to complex modifications in scenes, objects, and actions in natural images. Our method accurately interprets user intent and retrieves the desired images.
While our method yields promising results, it has certain limitations. As shown in Figure 4, although our method can retrieve images that align well with most user requirements, it sometimes struggles to capture the full semantics of complex multimodal queries, such as understanding abstract concepts like ‘simple print’ (1st row of Figure 4(a)), differentiating subtle differences like dog breeds (1st row of Figure 4(b)), or determining relative object positioning (2nd row of Figure 4(b)). These limitations indicate a need for mechanisms to enhance our method’s ability to interpret nuanced queries.
@@ 5 Conclusion
We have presented a training-free approach for ZS-CIR. By leveraging pretrained VLMs and MLLMs, our approach integrates information from different modalities via a weighted average. This results in an easy approach to construct the query representation and allows for the incorporation of textual information into the similarity computation. Our approach performs CIR without any training and hence has the flexibility to use different VLMs and MLLMs depending on the datasets and applications. Experiments and analyses on two datasets demonstrate the effectiveness of our method. Future work will focus on enhancing our method’s ability to interpret nuanced queries.
{credits}
5.0.1 Acknowledgements
This work was supported in part by the National Science and Technology Council, Taiwan under Grant NSTC 113-2640-E-110-002, 112-2221-E-110-047-MY3, and 112-2634-F-006-002.